OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!! Write the standard form of the equation of the line though the pair of points (-4,-1)and (-8,4). Write the equation using only integer coefficients...The equation is __x+__y=__

OpenStudy (anonymous):

y2-y1 / x2-x1 the equation is : 4-(-1) / -8-(-4) = 5/-4

OpenStudy (anonymous):

since the slope is 5/-4 then use : y+1 = 5/-4x +4 = Y = 5x - 1